Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Thursday, December 03, 1998

The following is a brief summary of the environmental monitoring activities that were completed on December 03, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

12/03/98 (Thursday)

Segment 1 Report for 12/3/98: Electrical work on the office interior at Emidio Pump Station continued as well as calibration of the pumps. Conduit was placed through the Hwy. 166 bore (EMP 1.18). Welding, x-ray and coating continued at the Stream Crossing #4 bore (EMP 9.03). Installation of the bore equipment continued at the Stream Crossing #9 bore (EMP 15.35). On Grapevine Rd., welding, x-ray and coating continued from 15.15.82 to 15.69. At Grapevine Pump Station electrical work continued. On Tejon Bench, backfill took place from EMP 16.89 to 16.85. A segment of pipe was tied-in at EMP 18.09. On Lebec Rd. a segment of pipe was tied in at EMP 21.53. Hydrostatic testing has been delayed due to lack of available water necessary to fill the pipes.

Segment 2B (Castaic to Sylmar): Report for 12-3-98. Electrical crews returned to Rye Cyn to complete work on the valve box. At the Santa Clara river, crews began a general clean up of the job site and removed sand bags from the rivers edge. Fiber optic crews pulled string through the duct between Pico Cy.. Rd and Calgrove Blvd. Welding crews worked to fabricate pipe at Weldon Cyn while trenching crews completed the tie in 700' south of Weldon Cyn Rd. Meanwhile, pipe crews laid pipe in the ditch to EMP 71.71 while welders worked to EMP 71.8. Coating and x-ray crews progressed just south of the welding crews while slurry and fiber optic crews progressed with shading the pipe near EMP 72.34.

Inspection Summary

In the past 24 hours, 3 inspection event(s) were reported for 3 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page